### v1

# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

Abstract
Multiple instance learning (MIL) has become a standard paradigm for weakly supervised classification of whole slide images (WSI). However, this paradigm relies on the use of a large number of labelled WSIs for training. The lack of training data and the presence of rare diseases present significant challenges for these methods. Prompt tuning combined with the pre-trained Vision-Language models (VLMs) is an effective solution to the Few-shot Weakly Supervised WSI classification (FSWC) tasks. Nevertheless, applying prompt tuning methods designed for natural images to WSIs presents three significant challenges: 1) These methods fail to fully leverage the prior knowledge from the VLM’s text modality; 2) They overlook the essential multi-scale and contextual information in WSIs, leading to suboptimal results; and 3) They lack exploration of instance aggregation methods. To address these problems, we propose a Multi-Scale and Context-focused Prompt Tuning (MSCPT) method for FSWC tasks. Specifically, MSCPT employs the frozen large language model to generate pathological visual language prior knowledge at multi-scale, guiding hierarchical prompt tuning. Additionally, we design a graph prompt tuning module to learn essential contextual information within WSI, and finally, a non-parametric cross-guided instance aggregation module has been introduced to get the WSI-level features. Based on two VLMs, extensive experiments and visualizations on three datasets demonstrated the powerful performance of our MSCPT.
Index Terms:
whole slide image classification, prompt tuning, few-shot learning, multimodal.
I Introduction
Figure 1: Motivation of our MSCPT. (a) Traditional MIL-based methods mainly focus on instance aggregation and require a large amount of training data. (b) Prompt tuning methods for natural images incorporate a set of trainable parameters into the input space for training, enabling pre-trained VLMs to be applied to downstream tasks. However, those methods are only suitable for single images and are no longer adequate for WSI-level tasks due to the enormous size of WSIs. (c) MSCPT leverages pathological visual descriptions combined with multimodal hierarchical prompt tuning to explore the potential of VLMs. For simplicity, we only depicted the data flow diagram for a single scale.
Developing automated analysis frameworks using Whole Slide Images (WSIs) is crucial in clinical practice [1, 2, 3, 4], as WSIs are widely regarded as the “gold standard” for cancer diagnosis, typing, staging, and prognosis analysis [5, 6]. Given the enormous size of WSIs (roughly 40,000 × 40,000), multiple instance learning (MIL) [7] has become the dominant method. As shown in Fig. 1 a, traditional MIL-based methods typically follow a four-step paradigm: patch cutting, feature extraction, feature aggregation, and classification. Most MIL-based methods are conducted under weak supervision at the bag level, as creating instance-level labels is quite labor-intensive [8, 9]. This weak supervision paradigm has led to a problem: a large number of WSIs are required to train an effective model [10, 11]. In clinical practice, patient privacy concerns, rare diseases, and difficulty preparing pathology slides make accumulating a large number of WSIs very challenging [12, 13].
Vision-Language models (VLMs) have shown excellent generalization ability to downstream tasks [14, 15, 16, 17, 18, 19, 20]. Recently, researchers have proposed specialized VLMs for analyzing pathological images, including MI-Zero [21], PLIP [22], and Conch [23]. These VLMs, extensively pre-trained on abundant image-text pairs, contain significant prior knowledge. If the prior knowledge of VLMs can be fully exploited with a few training samples, it can partially alleviate the data scarcity problem in WSI classification tasks. Therefore, we aim to explore a novel “data-efficient” method based on the VLMs to improve the model’s performance on the Few-shot Weakly Supervised WSI Classification (FSWC) [8] task.
Nevertheless, there is a gap between generally pre-trained VLMs and specific downstream tasks. Under the few-shot scenarios, researchers often employ prompt tuning to bridge this gap with the help of a few training samples [14, 24, 25, 18]. As shown in Fig. 1 b, prompt tuning aims to learn a set of trainable continuous vectors and incorporate these vectors into the input space for training, effectively adapting the fixed pre-trained VLMs for specific downstream tasks.
However, existing prompt tuning methods for natural images (such as CoOp [25], CoCoOp [24], and MetaPrompt [26]) are only effective for single images (i.e., patch-level). Since each WSI typically contains tens of thousands of patches, these methods are ineffective for WSI-level tasks. Also, studies indicate that the multi-scale information [27] and the contextual information [28] in WSIs play a significant role in cancer analysis, but those methods fail to capture this crucial information. Additionally, in training VLMs, the image-text pairs contain more than just information about the category. They also include more details about the image, such as contextual properties of the object [14] and descriptions of the cellular microenvironment [22, 23]. However, existing prompt tuning methods have primarily focused on image category information, without emphasizing a detailed image content analysis, which has left the full potential of pathological VLMs underexplored.
To address the aforementioned issue, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) for WSI classification in weakly supervised and few-shot scenarios. Our framework fully leverages the characteristic of VLM training with image-text pairs at dual magnification scales: 1) At low magnification, we provide the VLM with pathological visual descriptions at the tissue level (such as the infiltration between tumor tissue and other normal tissues); 2) At high magnification, pathological visual descriptions at the cellular level (such as cell morphology, nuclear changes, and the formation of various organelles) are provided to the VLM. These pathological visual descriptions at multi-scale can help VLM identify regions that are helpful for cancer analysis and achieve optimal results even with limited training samples.
As illustrated in Fig. 1 c, the core idea behind developing MSCPT is to incorporate prior knowledge at the tissue and cellular scales into the WSI-level tasks. Specifically, we first use a frozen large language model (LLM) to generate multi-scale pathological visual descriptions, leveraging them as prior knowledge.
Secondly, we design a Multi-scale Hierarchical Prompt Tuning (MHPT) module to combine pathological visual descriptions from multi-scale hierarchically to enhance prompt effectiveness. Inspired by Metaprompt [26], a dual-path asymmetric framework is adopted, asymmetrically freezing the image encoder and text encoder at different scales for prompt tuning. This asymmetric framework enables us to freeze half of the encoder to reduce the number of trainable parameters. Specifically, MHPT contains low-level and high-level prompts for both low and high-magnification visual descriptions, as well as global trainable prompts. The MHPT module employs the transformer layers in the text encoder to effectively learn the interactions among three distinct prompts.
Furthermore, the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module is introduced to extract contextual information. Precisely, we do not follow previous approaches [29, 30] of using patch positions or patch feature similarity to construct graph neural networks (GNNs). We propose to use the similarity between patches and pathological visual descriptions as the basis for building GNNs. We believe that using image-text pairs to build GNNs is more effective for capturing global features than methods relying on patch positions and image feature similarity, and corresponding ablation experiments confirm this hypothesis.
Finally, impressed by the powerful zero-shot capabilities of VLMs [22, 23, 21], we fully leverage the similarity between patches and pathological visual descriptions to aggregate instances. The Non-Parametric Cross-Guided Pooling (NPCGP) module, utilizing the Top-K algorithm for instance aggregation, is introduced to further reduce the risk of overfitting in few-shot scenarios. Overall, our contributions are summarised as follows:
  1. 1.
    MSCPT demonstrates that high-level concepts from pathological descriptions combined with low-level image representations can enhance few-shot weakly supervised WSI classification.
  2. 2.
    MSCPT achieves excellent performance by introducing only a limited number of trainable parameters ($\sim$0.9of the pre-trained VLM). Additionally, MSCPT is applicable to fine-tune any VLMs for WSI-level tasks.
  3. 3.
    Extensive experiments and visualizations on three datasets and two VLMs have confirmed that MSCPT’s performance is state-of-the-art in few-shot scenarios, surpassing other traditional MIL-based and prompt tuning methods.
Figure 2: We develop MSCPT based on the dual-path asymmetric framework, which inputs patches and pathological visual descriptions from multi-scale to different encoders. MSCPT utilizes a large language model to generate multi-scale pathological visual descriptions. These descriptions are combined using Multi-scale Hierarchical Prompt Tuning (MHPT) to integrate information across multiple scales. Then Image-text similarity-based Graph Prompt Tuning (ISGPT) is employed to learn context information at each scale. Finally, Non-Parametric Cross-Guided Pooling (NPCGP) aggregates instances guided by pathological visual descriptions to achieve the final Whole Slide Image classification result.
II Related Work
II-A Multiple Instance Learning in Whole Slide Images
Due to the high resolution of Whole Slide Images (WSIs) and the challenges of detailed labelling, weakly supervised methods based on Multiple Instance Learning (MIL) have emerged as the mainstream for WSI analysis. The MIL-based methods treat a WSI as a bag and all patches as instances, considering a bag positive if it contains at least one positive instance. Within the MIL framework, an aggregation step is required to aggregate all instances into bag features. The most primitive aggregation methods are non-parametric mean pooling and max pooling. However, since disease-related instances are a small fraction [31], those non-parametric aggregation methods treated all instances equally, causing useful information to be overwhelmed by irrelevant data. Subsequently, some attention-based methods (such as ABMIL [32], DSMIL [33] and CLAM [31]) were introduced, assigning different weights to each instance and aggregating them based on the weights. Furthermore, MIL methods based on Graph Neural Networks (GNNs) [29, 30] and Transformers [1, 34] had also been proposed to capture both local and global contextual information of WSIs. Those methods have shown significant improvements in recent years. Still, the cost of enhancing model performance is the increase in parameters, requiring a large amount of data to train a well-performing model. In many cases, training data faces a scarcity issue. Therefore, this paper proposes MSCPT, which leverages Vision-Language models combined with pathological descriptions from Large language models to enhance the performance in few-shot scenarios.
II-B Vision-Language Models
Vision-Language models (VLMs) are rapidly developing in various fields. During training, VLMs use contrastive learning to reduce distances between paired image-text pairs and increase distances between unpaired ones. CLIP [14] collected over 400M image-text pairs from the internet and used contrastive learning to align them, resulting in compatibility across various tasks. Compared to natural images, gathering pairs of pathological images and corresponding descriptions is challenging. To address this issue, MI-Zero [21] first pretrains image and text encoders using unpaired data, and then aligns them in a common space using 33,480 pairs of pathological image-text pairs. Huang et al. gathered over 450K pathological image-text pairs from Twitter and LAION [35] and developed PLIP [22]. Lu et al. trained Conch [23] on over 1.17M pathological image-text pairs, and it performs well on downstream tasks. Pretrained VLMs have significant potential, but effective methods to leverage them for WSI-level tasks are lacking. In this paper, we propose using pathological visual descriptions as prior knowledge to unleash the potential of VLMs.
II-C Prompt Tuning in Vision-Language Models
Prompt tuning has demonstrated remarkable efficiency and effectiveness, whether in text or multimodal [25, 24, 18]. CLIP demonstrated remarkable zero-shot performance with hand-crafted prompts, but the results can vary significantly depending on the prompt used due to their sensitivity to changes. Therefore, CoOp [25] and CoCoOp [24] proposed that the model itself should determine the choice of prompts. Khattak et al. argued that optimizing prompt tuning within a single branch is not ideal. They introduced MaPLe [18] as a solution to enhance the alignment between visual and language representations. Regrettably, these innovative methods are highly applicable to natural images but do not consider the enormous size of WSIs and the crucial multi-scale and contextual information needed for WSI analysis.
To our knowledge, Qu et al. have conducted research TOP [8] on the fine-tuning of CLIP for FSWC tasks. Shi et al. also proposed ViLa-MIL [36] based on CLIP, which helps with WSI classification by introducing multi-scale language prior knowledge. These two studies are exceptional, pushing the boundaries of VLM capabilities and boosting model performance in few-shot scenarios. However, these methods are all based on CLIP and do not investigate the performance of models on pathological VLMs. Moreover, due to the large number of patches in a WSI, they have to focus solely on the text and neglect visual prompt tuning. Additionally, they do not consider the crucial contextual information in WSI. Although ViLa-MIL takes into account multi-scale information, it merely integrates information using a late fusion approach without fully exploring the interactions between these scales.
We validated our proposed MSCPT on both general VLM (i.e., CLIP) and pathology-specific VLM (i.e., PLIP). By utilizing the zero-shot capability of VLM to initially select a subset of patches closely related to cancer, we then conducted visual prompt tuning on these patches. Additionally, we adopted an intermediate fusion approach to integrate pathology prior knowledge at multi-scale, leveraging the transformer layers to hierarchically learn the relationships between them. Ultimately, we also utilized image-text similarity to construct GNNs to capture contextual information within the WSI.
III Method
In this section, we introduce our few-shot weakly-supervised WSI classification model, named Multi-scale and Context-focused Prompt Tuning (MSCPT), as illustrated in Fig.  2. MSCPT utilizes a dual-path asymmetric structure as its foundation while conducting hierarchical prompt tuning on both textual and visual modalities.
III-A Problem Formulation
Given a dataset $X=\left\{X_{1},X_{2},...,X_{N}\right\}$ consisting of $N$ WSIs, each WSI is cropped into non-overlapping small patches, named instances. All instances belonging to the same WSI collectively form a bag. In weakly-supervised WSI tasks, only the labels of bags are known. The labels of the bags $Y_{i}\in\left\{0,1\right\},\ i=\{1,2,...N\}$ and the label of each instance $\{y_{i,j},j=1,2,\ldots M_{i}\}$ have the following relationship:
$Y_{i}=\left\{\begin{array}[]{cc}0,&\quad\text{if}\sum_{j}y_{i,j}=0,\\ 1,&\quad\text{else}.\end{array}\right.$ (1)
III-B Review of CLIP and Patch Selection
III-B1 Review of CLIP
CLIP [14] adopts a two-tower structure, including an image encoder and a text encoder. The image encoder $\mathcal{F}_{img}$ can be either a ResNet [37] or ViT [38], which is used to transform images into visual embeddings. The text encoder $\mathcal{F}_{text}$ takes a series of words as input and outputs textual embeddings. During the training process, CLIP utilizes a contrastive loss to learn a joint embedding space for the two modalities. During inference, we assume $\boldsymbol{x}$ is the visual embedding, and $\left\{\boldsymbol{w_{i}}\right\}_{i=1}^{K}$ is a series of textual embeddings generated by $\mathcal{F}_{text}$. Each $\boldsymbol{w_{i}}$ corresponds to prompt (such as “an image of {class name}) embedding for a specific image category. Therefore, the predicted probabilities can be obtained by calculating the cosine similarity between $\boldsymbol{x}$ and $\boldsymbol{w_{i}}$:
$p(y=i\mid\boldsymbol{x})=\frac{exp(cos(\boldsymbol{x},\boldsymbol{w_{i}})/\tau)}{{\textstyle\sum_{j=1}^{K}exp(cos(\boldsymbol{x},\boldsymbol{w_{j}})/\tau)}},$ (2)
where $\tau$ is the temperature coefficient, $cos(\cdot,\cdot)$ represents the cosine similarity, and $K$ is the number of categories.
III-B2 Patch Selection
Due to the high resolution of WSIs, dividing them into non-overlapping patches will result in a large number of patches. However, research has shown that only a few patches contain crucial information [31]. By preliminarily identifying patches closely linked to cancer analysis, we can notably diminish the computational resources demanded by visual prompt tuning. The powerful zero-shot ability of the VLMs allows for the initial screening of cancer-related patches.
Specifically, we utilize $\mathcal{F}_{img}$ to extract visual embeddings from patches while leveraging $\mathcal{F}_{text}$ to extract textual embeddings from the category prompts. Following this, the similarities between patches and prompts are computed. Then, the top $n$ patches with the highest similarity scores are selected for each category. To enhance the robustness of patch selection, we generated 50 sets of manual category templates and averaged their embeddings following  [21]. For a WSI $X_{i}$, we choose patches $\left\{x^{l}_{i,j},j=1,2,...,n_{l}\right\}$ at low magnification. Due to our unique architecture, we solely perform patch selection and visual prompt tuning at low magnification.
III-C Multi-scale Visual Descriptions Construction
In this part, we aim to generate pathological visual descriptions as pathological language prior knowledge to guide the hierarchical prompt tuning and instances aggregation. To reduce manual workload, large language models (LLMs) are employed to generate descriptions related to different diseases. That is, we enter the question “We are studying Cancer Category. Please list $\boldsymbol{C^{l}}$ visual descriptions at 5×5$\times$5 × magnification and $\boldsymbol{C^{h}}$ visual descriptions at 20×20$\times$20 × magnification observed in H&E-stained histological images of Cancer Sub-category.” into the LLM. And then we can get the multi-scale visual description sets $T^{low}=\{T^{low}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{l}\}$ and $T^{high}=\{T^{high}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{h}\}$. $K$ represents the number of WSI categories, and $C^{l}$ and $C^{h}$ denote the counts of low-level and high-level descriptions, respectively.
III-D Multi-scale Hierarchical Prompt Tuning
Inspired by MetaPrompt [26], a unique dual-path asymmetric framework is employed for multimodal hierarchical prompt tuning, as shown in the left of Fig. 2. Freezing two of all encoders helps reduce the trainable parameters and alleviates overfitting in few-shot scenarios. Compared to previous works where their encoders process the same inputs, our method adopts a unique strategy: the prompted and frozen encoders take entirely different inputs. Considering the immense size of WSIs and the substantial computational and storage resource requirements for visual prompt tuning, we only conducted visual prompt tuning at the low level.
Rather than modifying the visual prompts tuning method from Metaprompt, our emphasis is placed on the text modality. Specifically, the low-level pathological visual descriptions $T^{low}$ are sent into the frozen low-level text encoder $\mathcal{F}^{low}_{text}$, while the high-level pathological visual descriptions $T^{high}$ are sent into the prompted hierarchical high-level text encoder $\mathcal{F}^{high}_{text}$. Simultaneously, patches are also fed into the corresponding encoders. We wish to integrate different information contained in $T^{low}$ and $T^{high}$, which can help improve the multi-scale information processing capability of MSCPT. To achieve this purpose, we propose Multi-scale Hierarchical Prompt Turning (MHPT) module. The core component of MHPT, prompted hierarchical high-level text encoder, has been drawn in Fig. 3.
Figure 3: Details of the Prompted Hierarchical High-Level Text Encoder. The multi-scale Hierarchical Prompt Turning (MHPT) module utilizes the transformer layer to integrate pathological visual descriptions from different scales.
III-D1 Multi-scale Prompts Construction
For each layer of $\mathcal{F}^{low}_{text}$, we introduce a learnable vector called global prompts $p_{glob}$ to learn and integrate information from high-level text prompts $p_{high}$ and low-level text prompts $p_{low}$. As an example, consider the construction of multi-scale prompts for a high-level pathological visual description $T^{high}_{k,c}$. After tokenization and embedding, $T^{high}_{k,c}$ is transformed into $p^{high}_{0}$. And then the low-level text prompts $p_{low}$ are obtained based on $T^{low}$. More specifically, a set of descriptions $T^{low}$ gets fed into the frozen $\mathcal{F}^{low}_{text}$, and the last token of each transformer layer gets extracted. These tokens are then fed into a prompt generator $g$, formulated as:
$p_{low,i}^{l}=g\left(d_{low}^{l}\right),$ (3)
where $d_{low}^{l}$ is the last token of $T^{low}_{k,i}$ at the $l$-th layer, generator $g$ is a basic multilayer perceptron to align vectors of different scales into a common embedding space. Then, these tokens get concatenated to obtain low-level text prompts $p_{low}^{l}$.
III-D2 Hierarchical Prompt Tuning
After obtaining the three prompts, to capture more complex associations between pathological visual descriptions at multi-scale, hierarchical prompt tuning is performed on $\mathcal{F}^{high}_{text}$, which can be expressed as:
[…]
where $C^{i}$ and $E^{i}$ represent the class token $\left[CLS\right]$ and the last token $\left[EOT\right]$ of the $i$-th transformer layer $T^{i}$, and $L$ signifies the number of transformer layers. Lastly, by projecting the last token of the last transformer layer through the textual projection head $TextProj$ into the joint embedding space, the final textual representation $z^{high}_{k,c}$ for $T^{high}_{k,c}$ is obtained:
$z^{high}_{k,c}=Textproj(E^{L}).$ (5)
III-E Image-text Similarity-based Graph Prompt Tuning
Some studies have shown that the interactions between different areas of WSI and their structural information play a crucial role in cancer analysis [28]. However, the current prompt tuning methods are unable to capture this information. To address this, we propose Image-text Similarity-based Graph Prompt Tuning (ISGPT) module. More specifically, we deviate from conventional methods that utilize patch coordinates or patch feature similarity in constructing graph neural networks (GNNs) [29, 30]. Our innovative approach involves utilizing the similarity between patches and pathological visual descriptions as the foundation for developing GNNs. We treat the patches as nodes and aim to construct the adjacency matrix $\boldsymbol{A}$ by calculating the semantic similarity $\boldsymbol{S}$ between the patch embeddings and description embeddings. Specifically, after patches and descriptions have passed through the encoders from Section III-D, patch embeddings $\boldsymbol{P}\in\mathbb{R}^{M\times d}$ and description embeddings $\boldsymbol{Z}\in\mathbb{R}^{KC\times d}$ are obtained, respectively. The formula for semantic similarity $\boldsymbol{S}\in\mathbb{R}^{M\times KC}$ is:
$s_{i,j}=\frac{exp(cos(\boldsymbol{P_{i}},\boldsymbol{Z_{j}})/\tau)}{{\textstyle\sum_{m=1}^{K\times C}exp(cos(\boldsymbol{P_{i}},\boldsymbol{Z_{m}})/\tau)}},$ (6)
where $\tau$ is the temperature coefficient, $cos(\cdot,\cdot)$ represents the cosine similarity. $K$ represents the number of WSI categories and $d$ is the embedding dimensionality. $C$ and $M$ denote the number of pathological descriptions and patches at a given scale, respectively. Subsequently, the calculation formula for the adjacency matrix $\boldsymbol{A}\in\mathbb{R}^{M\times M}$ is written as:
$a_{i,j}=\frac{exp(cos(\boldsymbol{S_{i}},\boldsymbol{S_{j}})/\tau)}{{\textstyle\sum_{m=1}^{M}exp(cos(\boldsymbol{S_{i}},\boldsymbol{S_{m}})/\tau)}},$ (7)
where $\boldsymbol{S_{i}}\in\mathbb{R}^{KC}$ represents the semantic similarity between $i$-th patch embeddings and all description embeddings(i.e., the $i$-th row of $\boldsymbol{S}$). We avoid constructing $\boldsymbol{A}$ based on patch coordinates or patch feature similarity, as this approach might overlook fewer but significant patches when focusing only on Euclidean distance or patch feature similarity. Subsequent experimental results have demonstrated the superior performance of our method for constructing $\boldsymbol{A}$. We choose Graph Convolutional Network (GCN) [39] as the graph learning model. The definition of the GCN operation in the $l$-th GCN layer is as follows:
$\mathcal{F}_{GCN}\left(\boldsymbol{A},\boldsymbol{H}^{\left(l\right)}\right)$ $=\sigma(\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\tilde{\boldsymbol{A}}\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\boldsymbol{H^{\left(l\right)}}\boldsymbol{W}^{\left(l\right)}).$ (8)
Here $\tilde{\boldsymbol{A}}=\boldsymbol{A}+\boldsymbol{I}$, $\boldsymbol{I}$ is the identity matrix and $\sigma\left(\cdot\right)$ denotes an activation function. $\tilde{D}_{i,i}={\textstyle\sum_{j}\tilde{A}_{i,j}}$, $\boldsymbol{W}^{\left(l\right)}$ is layer-specific trainable weight matrix. $\boldsymbol{H}^{\left(l\right)}\in\mathbb{R}^{M\times d}$ is the input embeddings of all nodes. Therefore, the patch embeddings after graph prompt tuning at both high and low scales are represented as:
$\tilde{\boldsymbol{P}}^{high}$ $=\mathcal{F}^{high}_{GCN}(\boldsymbol{A}^{high},\boldsymbol{P}^{high}),$ (9)
$\tilde{\boldsymbol{P}}^{low}$ $=\mathcal{F}^{low}_{GCN}(\boldsymbol{A}^{low},\boldsymbol{P^{low}}).$ (10)
III-F Non-Parametric Cross-Guided Pooling
Impressed by the powerful zero-shot capability of pre-trained VLMs, the possibility of employing a similar non-parametric approach for instance aggregation was pondered. We propose Non-Parametric Cross-Guided Pooling (NPCGP) to aggregate instance into bag features. In NPCGP, we compute semantic similarities between the patch embeddings $\tilde{\textbf{P}}$ post graph tuning and pathological visual description embeddings $\boldsymbol{Z}$ at both the same and across scales. The reason for calculating similarities both within the same and across scales is our concern that the pathological visual descriptions provided by LLM may contain scale-related inaccuracies. Hence, this procedure serves to bolster the robustness of feature aggregation strategies. Lastly, the bag-level unnormalized probability distribution $Logits$ is obtained through the topK max-pooling operator $h_{topK}$:
$Logits^{high}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right),\end{aligned}$ (11)
$Logits^{low}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right),\end{aligned}$ (12)
[…]
Following previous work [26], we use cross-entropy loss to optimize the three distributions $Logits^{overall}$, $Logits^{high}$, and $Logits^{low}$, but only $Logits^{overall}$ was used during model inference.
TABLE I: Cancer sub-typing results on TCGA-NSCLC, TCGA-BRCA, and TCGA-RCC. The highest performance is in bold, and the second-best performance is underlined. We provided mean and standard deviation results under five random seeds.
{tblr}
colsep=2pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, row16 = bg=lightgray, row30 = bg=lightgray, cell11 = c=2,r=2l, cell13 = r=2, cell15 = c=3, cell19 = c=3, cell113 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, hline1,17,31 = -0.08em, hline2 = 5-7,9-11,13-15, hline3 = -, Methods & Trainable
Param TCGA-NSCLC TCGA-BRCA TCGA-RCC
AUC F1 ACC AUC F1 ACC AUC F1 ACC
CLIP ImageNet Pretrained Max-pooling 197K 63.80±6.84 60.40±4.76 60.70±4.75 60.42±4.35 56.40±3.58 68.552±6.54 84.51±3.21 65.83±2.72 69.26±2.33
Mean-pooling 197K 69.53±4.74 63.76±5.77 64.69±4.31 66.64±2.41 60.70±2.78 71.73±3.59 93.31±0.66 78.64±0.74 81.29±1.08
ABMIL [32] 461K 66.95±4.31 62.60±3.75 62.96±3.65 67.92±3.90 61.72±3.60 72.77±3.15 93.41±1.41 79.80±1.56 82.47±1.46
CLAM-SB [31] 660K 67.49±5.94 62.86±4.19 63.51±4.19 67.80±5.14 60.51±5.07 72.46±4.36 93.85±1.52 79.87±3.17 83.21±2.67
CLAM-MB [31] 660K 69.65±3.61 64.52±3.22 65.14±2.69 67.98±4.86 60.68±6.47 74.09±3.52 93.59±1.16 78.72±2.18 8103±2.06
TransMIL [1] 2.54M 64.82±8.01 59.17±10.87 62.00±5.18 65.31±6.02 57.72±2.48 68.12±4.11 94.17±1.23 79.63±1.52 81.86±1.41
DSMIL [33] 462K 66.00±9.23 63.87±7.00 64.11±6.65 66.18±10.08 59.35±8.01 67.52±11.56 91.53±5.17 78.38±6.56 80.69±6.47
RRT-MIL [40] 2.63M 66.47±6.73 62.10±6.17 63.20±5.24 66.33±4.30 61.14±5.93 71.21±8.94 93.89±1.91 81.04±2.11 83.30±2.24
CoOp [25] 337K 69.06±4.06 63.87±3.77 64.27±3.55 68.86±3.45 61.64±2.40 72.10±3.22 94.18±1.72 79.88±2.40 82.15±1.96
CoCoOp [24] 370K 64.37±2.28 60.95±1.55 61.37±1.36 66.50±3.02 59.64±2.90 71.07±4.93 85.68±2.66 67.72±3.49 71.00±2.90
Metaprompt [26] 360K 75.94±3.01 70.35±3.09 70.41±3.09 69.12±4.12 63.39±4.28 74.65±7.20 94.18±1.56 80.03±2.06 82.52±2.15
TOP [8] 2.11M 73.56±3.14 68.19±1.22 68.77±2.53 69.75±4.66 61.32±6.12 71.68±2.56 93.56±1.22 79.66±1.97 80.79±1.05
ViLa-MIL [36] 2.77M 74.85±7.62 68.74±5.86 68.87±5.97 70.13±3.86 62.04±2.28 71.93±2.31 93.34±1.49 79.40±1.13 81.81±0.92
[…]
PLIP Pathology Pretrained Max-pooling 197K 71.78±4.13 66.40±3.51 66.66±3.42 66.66±2.36 60.32±2.24 71.57±4.83 95.18±0.63 81.63±0.92 84.30±1.30
Mean-pooling 197K 70.55±6.64 65.32±5.60 65.50±5.55 71.62±2.41 64.62±2.96 74.45±2.49 94.75±0.51 82.22±0.67 85.24±0.80
ABMIL [32] 461K 78.54±4.29 72.06±3.79 72.12±3.78 72.18±1.28 64.49±1.74 74.63±1.31 96.51±0.63 85.66±1.97 87.94±1.92
CLAM-SB [31] 660K 80.56±4.57 73.15±4.05 73.27±3.97 73.49±2.12 65.22±2.61 75.05±3.88 96.41±0.36 84.71±1.60 87.25±1.34
CLAM-MB [31] 660K 80.68±3.63 73.15±3.00 73.32±2.83 74.33±1.76 66.11±1.94 76.11±2.03 96.58±0.59 85.20±0.83 87.85±0.79
TransMIL [1] 2.54M 73.40±10.33 66.92±7.94 67.21±7.63 70.52±2.45 62.06±1.67 70.14±2.77 96.35±0.54 83.70±0.80 86.33±0.46
DSMIL [33] 462K 77.75±7.22 72.84±6.31 73.08±6.00 70.14±4.11 63.01±2.78 71.48±5.37 93.01±6.05 79.58±9.16 82.87±7.32
RRT-MIL [40] 2.63M 76.30±10.01 70.86±7.47 71.01±7.44 72.77±2.20 65.74±2.34 74.38±4.01 96.09±1.06 83.94±2.05 86.56±2.28
CoOp [25] 337K 77.92±5.48 71.58±4.74 71.63±4.75 73.77±2.83 64.88±1.26 74.14±3.38 95.76±0.80 83.23±2.07 85.90±1.63
CoCoOp [24] 370K 72.62±8.45 66.63±5.83 66.97±5.85 71.21±4.20 62.95±3.95 73.57±6.31 95.81±0.42 83.18±1.35 86.02±1.03
Metaprompt [26] 360K 78.31±5.66 72.03±4.60 71.86±4.61 73.98±2.15 65.50±2.05 75.56±4.58 95.75±0.48 83.52±1.46 86.62±1.43
TOP [8] 2.11M 78.91±3.79 72.33±4.89 72.91±4.61 74.06±2.66 65.17±2.16 76.51±1.79 95.06±0.51 82.86±1.35 86.14±0.98
ViLa-MIL [36] 2.77M 80.98±2.52 73.81±3.64 73.94±3.56 74.86±2.45 66.03±1.81 77.35±1.63 95.72±0.60 83.85±1.10 86.53±1.03
MSCPT(ours) 1.35M 84.29±3.97 76.39±5.69 76.54±5.49 75.55±5.25 67.46±2.43 79.14±2.63 96.94±0.36 87.01±1.51 89.28±1.22
IV Experimental Results
IV-A Experimental Settings
IV-A1 Datasets
To comprehensively assess the performance of our Multi-Scale and Context-focused Prompt Tuning (MSCPT), three real datasets from the Cancer Genome Atlas (TCGA) Data Portal were used: TCGA-NSCLC, TCGA-BRCA, and TCGA-RCC.
TCGA-NSCLC is a dataset of 1041 non-small cell lung cancer (NSCLC) WSIs, including 530 lung adenocarcinoma (LUAD) and 511 lung squamous cell carcinoma (LUSC) slides. 20% of the dataset (209 slides) is used for training, and the remaining 80% (832 slides) is used for testing.
TCGA-BRCA is a dataset comprising 1056 slides of breast invasive carcinoma (BRCA) WSIs. This dataset includes 845 slides of invasive ductal carcinoma (IDC) and 211 slides of invasive lobular carcinoma (ILC). 20% of them (223 slides) are randomly selected as the training set, and the remaining 80% (833 slides) are used as the testing set.
TCGA-RCC is a renal cell carcinoma (RCC) WSIs dataset containing 873 slides. Precisely, it consists of 121 slides of chromophobe renal cell carcinoma (CHRCC), 455 slides of clear-cell renal cell carcinoma (CCRCC), and 297 slides of papillary renal cell carcinoma (PRCC). Likewise, 20% of the dataset (175 slides) is randomly taken out for training, while 698 slides are reserved for testing.
IV-A2 Evaluation Metrics
For all datasets, we leverage Accuracy (ACC), Area Under Curve (AUC), and macro F1-score (F1) to evaluate model performance. To reduce the impact of data split on model evaluation, we follow ViLa-MIL [36] and employ five fixed seeds to perform five rounds of dataset splitting, model training, and testing. We report the mean and standard deviation of the metrics over five seeds.
IV-A3 Model Zoo
Thirteen influential approaches were employed for comparison, including traditional MIL-based methods: Mean pooling, Max pooling, ABMIL [32], CLAM [31], TransMIL [1], DSMIL [33] and RRT-MIL [40]; prompt tuning methods for natural images: CoOp [25], CoCoOp [24] and Metaprompt [26]; prompt tuning methods for WSIs: TOP [8] and ViLa-MIL[36]. Adapting to WSI-level tasks, we integrated an attention-based instance aggregation module [32] into the prompt tuning methods designed for natural images, such as CoOp, CoCoOp, and Metaprompt.
IV-A4 Implementation Details
Following CLAM [31], the original WSIs were initially processed using the Otsu thresholding algorithm to remove the background parts. Subsequently, the WSIs were segmented into multiple non-overlapping patches of $256\times 256$ pixels at 5×5$\times$5 × and 20×20$\times$20 × magnification levels. We applied to perform our MSCPT on CLIP [14] and PLIP [22], both of which use ViT-B/16 [38] as their visual tower. Apart from MSCPT, Metaprompt, and DSMIL, which utilized inputs of both 5×5$\times$5 × and 20×20$\times$20 × magnification patches, the remaining methods solely relied on 20×20$\times$20 × magnification patches as inputs.
For all methods, the Adam optimizer was employed with a learning rate of 1e-4, a weight decay of 1e-5 and batch size was set to 1. All methods were trained for a fixed number of epochs (100 for CLIP and 50 for PLIP) with early stop. We chose GPT-4 [41] to generate pathological visual descriptions, providing 10 low-level visual descriptions and 30 high-level visual descriptions for each category of WSIs (i.e., $C^{l}=10$ and $C^{h}=30$). For MSCPT and Metaprompt, we utilized the zero-shot capability of the VLMs to select 30 patches for each category at 5×5$\times$5 × magnification. The lengths of the global prompts $p_{glob}$ in both image and text encoder were uniformly set to 2. In this paper, unless explicitly stated otherwise, all experiments are conducted with 16 training samples per category. All the work was conducted using the PyTorch library on a workstation with eight NVIDIA A800 GPUs. All codes and details are released at https://github.com/Hanminghao/MSCPT.
IV-B Comparisons with State-of-the-Art
The experimental results under the 16-shot setting are displayed in Table I. We observed some intriguing insights, such as complex and parameter-heavy methods like TransMIL and RRT-MIL underperformed despite their strong performance with full data training. Conversely, less parameterized methods such as ABMIL and CLAM exhibited slightly better performance. This is because traditional MIL-based methods require a lot of WSIs for training and the more parameters they have, the more training data is needed. Furthermore, after adapting the prompt tuning methods designed for natural images (i.e., CoOp, CoCoOp, and Metaprompt) to tasks at the WSI level, these methods outperform traditional MIL-based methods when based on CLIP and achieve comparable performance when using PLIP. Relatively few parameters contribute to this result. Additionally, we found that Metaprompt outperforms CoOp across most metrics, thanks to its integration of visual prompt tuning and multi-scale information. This result motivates us to pursue visual prompt tuning and develop more effective multi-scale information integration modules. Despite prompt tuning methods designed for Few-shot Weakly Supervised WSI Classification tasks having a relatively higher number of parameters, they exhibit the best performance. This is because VLMs’ prior knowledge is effectively exploited under the guidance of visual descriptive text prompts, alleviating the demand for extensive training data.
Compared to other methods, our proposed MSCPT exhibits significant improvements in all evaluation metrics across the three datasets and two VLMs. Compared to the top-performing traditional MIL-based methods, MSCPT shows improvements of 0.3-13.0% in AUC, 2.0-12.3% in F1, and 1.5-11.6% in ACC across three datasets and two VLMs. Overall, MSCPT shows greater performance improvements when based on CLIP compared to PLIP. This is attributed to the specialized pre-training of PLIP on pathological images, enhancing its encoding capabilities for patches, thereby reducing the reliance on textual descriptions. Compared to the top-performing prompt tuning method suitable for natural images, MSCPT improved the AUC, F1, and ACC by 1.0-8.2%, 2.8-6.7%, and 1.6-6.9%.
TABLE II: Core components ablation experiment on the TCGA-NSCLC dataset based on PLIP.
MHPT ISGPT NPCGP TCGA-NSCLC (PLIP-based)
AUC F1 ACC
- - - 78.31±5.66 72.03±4.60 71.86±4.61
- - 80.57±4.17 73.62±2.41 73.77±2.48
- 82.75±5.73 75.48±4.70 75.53±4.72
- 81.92±5.03 74.96±4.68 75.10±4.64
84.29±3.97 76.39±5.69 76.54±5.49
TABLE III: Ablation experiment of different graph construction and training methods on the TCGA-RCC dataset based on CLIP.
Methods Trainable Param TCGA-RCC (CLIP-based)
AUC F1 ACC
GCN+KNN(Coord.) 1.35M 92.85±2.43 79.29±3.98 81.63±3.74
GCN+KNN(Feat.) 1.35M 93.92±2.66 80.46±4.33 82.41±4.26
GAT+Sim. 1.35M 93.14±1.78 80.82±4.25 82.41±3.85
GraphSAGE+Sim. 2.40M 93.60±2.72 80.40±3.89 82.66±3.69
GCN+Sim.(ours) 1.35M 95.04±1.31 82.59±2.14 85.62±2.14
Prompt tuning methods explicitly designed for WSI exhibit superior performance. This is attributed to their incorporation of priors into pre-trained VLMs and leveraging those priors to guide prompt tuning. Additionally, ViLa-MIL introduces multi-scale information compared to TOP, positioning it as the second-best overall performer. In comparison to ViLa-MIL, MSCPT shows improvements across all datasets, with AUC increasing by 0.9-5.7%, F1 by 2.2-6.2%, and ACC by 2.3-5.5%. This is because MSCPT performs prompt tuning both on the textual and visual modality. Furthermore, MSCPT takes into account both the multi-scale and contextual information of WSIs. Unlike the late fusion approach in ViLa-MIL, MSCPT employs an intermediate fusion method, leveraging the transformer layer and trainable global prompts to integrate pathological visual descriptions from both high and low levels.
IV-C Ablation Experiment
IV-C1 Effects of Each Component in MSCPT
To verify the effectiveness of three core components, ablation experiments were conducted on the TCGA-NSCLC dataset based on PLIP, the experimental results are presented in Table II. When all modules were removed, MSCPT regresses to the baseline (i.e., Metaprompt). All metrics showed significant improvement (2.2%-2.9%) after adding the Multi-scale Hierarchical Prompt Tuning (MHPT) module to the baseline. This is because the MHPT module utilizes transformer layers to integrate pathological visual descriptions across different scales, enhancing the model’s information aggregation capabilities. Building upon this, we introduced the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module, which led to improvements in all metrics (2.4%-2.7%). This demonstrated that utilizing ISGPT for contextual learning also enhances model performance, reaffirming the importance of contextual information for WSI analysis. When we added both the MHPT and Non-Parametric Cross-Guided Pooling (NPCGP) module to the baseline, in comparison to solely adding MHPT, the metrics improved by 1.7%-1.8%. This indicates that the NPCGP module, compared to attention-based pooling, is more effective in identifying important patches within the WSI, resulting in better instance aggregation results. When all modules work together, the baseline is transformed into MSCPT. MSCPT has shown improvements of 7.6% in AUC, 6.1% in F1, and 6.5% in ACC compared to the baseline.
TABLE IV: Ablation experiment of different instance aggregation methods on the TCGA-NSCLC dataset based on PLIP.
Methods TCGA-NSCLC (PLIP-based)
AUC F1 ACC
Mean Pooling 78.88±4.61 73.60±3.38 73.39±3.68
Max Pooling 82.63±4.58 75.68±3.51 75.89±3.54
Attention-based Pooling 82.75±5.73 75.48±4.70 75.53±4.72
NPCGP w/o cross-guidance 83.61±5.68 75.98±5.35 75.81±5.21
NPCGP(ours) 84.29±3.97 76.39±5.69 76.54±5.49
TABLE V: Results of different large language models on the TCGA-RCC dataset based on CLIP.
Methods TCGA-RCC (CLIP-based)
AUC F1 ACC
Gemini-1.5-pro [42] 94.14±1.79 81.61±2.69 83.67±3.63
Claude-3 [43] 93.97±2.35 80.61±4.05 82.72±3.59
Llama-3 [44] 94.63±1.42 82.36±1.81 84.38±2.08
GPT-3.5 [45] 94.52±2.16 82.09±3.07 84.27±3.61
GPT-4 [41] 95.04±1.31 83.78±2.19 85.62±2.14
Figure 4: Visualization of the original WSI, similarity score map for patch selection, heatmaps generated using MSCPT, attention heatmaps of the baseline (i.e., Metaprompt) and the best-performing traditional MIL-based method, selected high similarity patches and patches with highest similarity scores using MSCPT. The area surrounded by the red line in the original WSI is the tumor area.
Figure 5: Experiments on TCGA-BRCA with 16, 8, 4, 2, and 1-shot settings. (a) (b) are CLIP-based results, while (c) (d) are PLIP-based results.
IV-C2 Effects of Graph Construction
To validate the effectiveness of building adjacency matrices based on image-text similarity, we used K-Nearest Neighbor (KNN) to create adjacency matrices based on patch coordinates or visual features as referenced in studies [29, 30]. Additionally, we tested the effectiveness of GCN by comparing them with GAT [46] and GraphSAGE [47]. Experimental results on TCGA-RCC using CLIP are shown in Table III, where KNN(Coord.) and KNN(Feat.) refer to using KNN for constructing adjacency matrices based on patch coordinates and patch features, and Sim. signifies constructing adjacency matrices using image-text similarity. Switching to KNN to construct the adjacency matrix led to decreased performance across all metrics, whether based on coordinates or patch visual features. This decline may be attributed to the limited scope of connectivity in the adjacent matrix construction methods. Connecting only nearby patches based on coordinates restricts the GNN to the local context, while connecting visually similar patches based on their features may lack global information about interactions between different types of tissue organization. In contrast, our ISGPT module connects patches related to specific cancer types, overcoming local or visual similarity connection limitations and enabling a more comprehensive contextual understanding. When we replaced GCN with GAT or GraphSAGE, the model’s performance also experienced varying degrees of decline. We believe complex graph neural networks are unsuitable for few-shot scenarios.
IV-C3 Effects of Instance Aggregation
To validate the effectiveness of our instance aggregation method, we compared Non-Parametric Cross-Guided Pooling (NPCGP) with other aggregation methods (i.e., Mean Pooling, Max Pooling, and Attention-based Pooling). The experimental results based on PLIP on the TCGA-NSCLC are presented in Table IV. When we replaced NPCGP with other methods, the performance of the models decreased to varying degrees. This implies that our NPCGP can discern more impactful patches and aggregate them into bag features. The visualization results in Section IV-D support this point. We also conducted an ablation study on cross-guidance. Instead of computing cross-scale cosine similarity during feature aggregation, we only calculated cosine similarity between patch and description embeddings at the same scale. Removing cross-guidance led to a drop in performance across all metrics, likely because LLMs may produce descriptions with incorrect scales.
IV-C4 Effects of Large Language Models
To verify the impact of different LLMs on model performance, we compared the performance of MSCPT when using descriptions generated by different LLMs (i.e., Gemini-1.5-pro [42], Cluade-3 [43], Llama-3 [44], GPT-3.5 [45] and GPT-4 [41]). The results obtained using CLIP on TCGA-RCC are presented in Table V. When generating descriptions using Claude-3, MSCPT performs comparably to the baseline. However, MSCPT outperforms the baseline when using other LLMs. This demonstrates MSCPT’s robustness across different LLMs and highlights the benefit of accurate pathological visual descriptions for model performance. Providing a more accurate description helps improve model performance.
IV-D Visualization
As shown in Fig. 4, we have visualized a case of TCGA-RCC based on PLIP and a case of TCGA-BRCA based on CLIP. As depicted in Fig. 4a, during patch selection, CLIP assigned high similarity scores not only to tumor regions but also to non-tumor areas. This outcome arose because CLIP was not specifically designed for pathological images, resulting in a less-than-optimal zero-shot capability for this type of imagery. However, after prompt tuning using MSCPT, the model correctly assigned high scores to the actual tumor regions, while the regions that originally received high scores dropped to lower score ranges (red arrows in Fig. 4a). Meanwhile, CLAM-MB struggled to differentiate tumor and no-tumor. Similarly, Metaprompt assigned high attention weights to certain non-tumor tissues (red arrows in Fig. 4a).
During selecting patches using PLIP, the model could roughly identify tumor regions but also assigned high scores to a small number of non-tumor areas. However, this issue was mitigated with MSCPT (red arrows in Fig. 4b). While ABMIL could also determine instance importance, it tended to assign higher scores to certain non-tumor regions compared to MSCPT (yellow arrows in Fig. 4b). Due to PLIP’s improved ability to represent pathological images, Metaprompt produced visualization results comparable to MSCPT.
IV-E Results with Fewer Training Samples
To further validate MSCPT’s performance, we conducted experiments on TCGA-BRCA with 16, 8, 4, 2, and 1-shot settings. Based on the results in Table I, we selected several well-performing models (i.e., CLAM-MB, CoOp, Metaprompt, ViLa-MIL, and MSCPT) for these experiments. It is also worth noting that with limited training samples, sample selection significantly impacts model performance [8]. To address this, we conducted dataset splitting, model training, and testing using ten different seeds, excluding the two best and two worst results to calculate the average. Due to the sample imbalance in TCGA-BRCA, we just reported AUC and macro F1-score, as shown in Fig. 5. When using CLIP as the base model, MSCPT underperforms compared to CLAM-MB and Metaprompt in 1- and 2-shot settings, likely due to MSCPT’s larger parameter size and CLIP’s limited understanding of pathology descriptions. However, with 4 or more shots, MSCPT significantly outperforms other methods. Additionally, when using PLIP as the base model, MSCPT consistently performs better than any other method.
V Conclusion
In this paper, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) to solve the Few-shot Weakly-supervised WSI Classification (FSWC) task. MSCPT generates multi-scale pathological visual descriptions using GPT-4, guiding hierarchical prompt tuning and instance aggregation. Experiments on three WSI subtyping datasets and two Vision-Language models (VLMs) show that MSCPT achieved state-of-the-art results in FSWC tasks. Furthermore, MSCPT is applicable to fine-tune any VLMs for WSI-level tasks. However, we find that the model performance varies significantly on different datasets and VLMs. That is because the performance of model fine-tuning largely depends on the pre-trained VLM itself. We look forward to the emergence of more comprehensive and powerful pre-trained pathological VLMs, which will significantly promote the development of FSWC tasks and even all computational pathology.